(4L²−0.2L³)=0 find average production of labor marginal production 



1
Expert's answer
2022-02-04T07:48:05-0500

"4L^2-0.2L^3=0,"

"L^2(4-0.2L)=0,"

"L \not =0,"

"4-0.2L=0,"

"L=\frac 4{0.2},"

"L=20."